OpenStudy (anonymous):

Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve 24 - 3x = -27.

jagr2713 (jagr2713):

well if you subtract -27-24=-51 -3x=-51 then divided by3 x=17

OpenStudy (anonymous):

Thanks